# AC-LIO: Towards Asymptotic and Consistent Convergence in LiDAR-Inertial Odometry

### v1

@@ III-D Convergence Criteria
In order to exploit the statistical properties of the pointcloud, during the iterative state estimation, we select the mean value of the point-to-plane residual (abbreviated as APR above), as the criteria for evaluating the convergence of pointcloud registration. For simplification, assuming the pointcloud registration fully converges to the global optimum, considering the ranging accuracy of LiDAR and the construction of point-to-plane residual, we can roughly model the expectation of the APR, to provide a threshold reference for the convergence criteria.
Denote the fitted plane in pointcloud map as $S$ with its normal vector as $\mathbf{u}$, and the point-plane residual as $\mathbf{z}$. Then, the incident angle $\phi$ of point $\mathbf{p}$ with respect to plane $S$ can be expressed as:
$\begin{gathered}\phi=\arccos\dfrac{|\mathbf{u}\cdot\mathbf{p}|}{|\mathbf{u}|\cdot|\mathbf{p}|}\end{gathered}$ (11)
@@
When the laser beam is perpendicular to plane $S$, then the incident angle $\phi$ will be zero. Assuming that the ranging error $\mathbf{n_{p}}$ of laser point follows Gaussian distribution with zero mean and $\sigma$ variance, then the residual $\mathbf{z}_{\perp}$ will follow chi-squared distribution and can be expressed as follows:
@@
$\begin{gathered}\mathbf{p}^{L}=\mathbf{p}_{gt}^{L}+\mathbf{n_{p}},\;\mathbf{n_{p}}\sim\mathcal{N}(0,\sigma)\\ \mathbf{z}_{\perp}=|\mathbf{p}_{gt}^{L}-\mathbf{p}^{L}|\\ \frac{\mathbf{z}_{\perp}^{2}}{\sigma^{2}}\sim\chi^{2}(1),\;\mathbb{E}(\mathbf{z}_{\perp})=\sigma\end{gathered}$ (12)
@@
As the surrounding environment varies during ego-motion, the distribution of normal vector $\mathbf{n}$ of plane $S$ will also change. Eventually, the variation of the geometrical scene will be reflected in the distribution of the incident angle $\phi$. Assuming the incident angle is uniformly distributed over the domain $[0,\pi/2)$ , then the expected mean of the point-plane residual $\mathbf{z}$ can be expressed as:
$\begin{gathered}\mathbb{E}(\mathbf{z})=\frac{2}{\pi}\int_{0}^{\frac{\pi}{2}}\mathbb{E}(\mathbf{z}_{\perp})\cos{\phi}\;d\phi=\frac{2}{\pi}\sigma\end{gathered}$ (13)
@@
Therefore, assuming the pointcloud registration converges to the global optimum, and only considering the LiDAR ranging error, the expectation of APR should be ${2\sigma}/{\pi}$. If the mean value of current point-plane residual is lower than this criteria, then the last registration can be considered as sufficiently converged.
@@
According to the threshold of convergence criteria, we can easily determine the convergence degree of pointcloud registration, and control the backpropagation and the asymptotic distortion compensation. Noting that $\sigma_{t}=\mathbb{E}(\mathbf{z})$ and the selected threshold of convergence criteria as $\eta$, due to the factors such as dynamic objects and cumulative errors in pointcloud map, the threshold $\eta$ can be appropriately extended to the interval of $[\sigma_{t},2\sigma_{t}]$ in practice.
@@
In particular, we perform two conditional judgments in iterative state estimation: First, evaluate the previous LiDAR frame, if the APR is less than the threshold, it is considered to be sufficiently converged and we can continue. Then, check the convergence of current registration, if the APR is still less than the threshold, there is no need to back-propagate for small residual distortion; otherwise, perform the backpropagation and asymptotic compensation, in order to promote the consistent convergence in state estimation.
@@
The approach, on one hand, can ensure the backpropagation and asymptotic compensation only performed as the previous LiDAR frame converged well, thus enhancing the odometry robustness in long-time, large-scale scenario; on the other hand, can avoid lots of additional operations in unreliable or unnecessary conditions, further saving the system computational expense. The experiment results of our framework are shown in the next section.
@@ IV EXPERIMENT RESULTS
In this section, we qualitatively and quantitatively evaluate AC-LIO framework through a series of experiments with various public and self-collected datasets. To examine the accuracy improvement of our approach in long-time and large-scale scenarios, and the effect of convergence criteria on system robustness, we conduct extensive experiments on the odometry accuracy and the validity of convergence criteria.
IV-A Odometry Accuracy Evaluation
In the odometry accuracy experiments, we compare AC-LIO framework with current state-of-the-art LIO systems including FAST-LIO2 and Point-LIO. All algorithms are implemented in C++ and employed on an i9 CPU desktop computer with the ROS operating system in Ubuntu. These methods only utilize CPU for computation without the assistance of GPU parallel computing. Besides, in the odometry accuracy experiments, we maintain the threshold of convergence criteria in the extended interval, i.e., $\eta\in[\sigma_{t},2\sigma_{t}]$. A brief introduction of the public and self-collected handheld datasets is provided below:
@@ IV-A Odometry Accuracy Evaluation
TABLE I: Odometry Accuracy Evaluation - RMSE of ATE ($m$)
AC-LIO(Ours) FAST-LIO2 Point-LIO
Residence_5.2
0.415 0.505 1
Park_3.2
0.582 0.618
Subway_3.3
1.670 3.752
Mall_6.3
0.216 0.208 15.409
Garden_1005
0.773 1.053 0.917
Garden_1006
1.796 2.075 2.471
Garden_1008
1.152 1.577 1.590
  • 1
    –  means that the system severely diverged midway.
TABLE II: Odometry Accuracy Evaluation - End to end errors ($m$)
AC-LIO(Ours) FAST-LIO2 Point-LIO
Indoor_1
(~100m)
0.8929 1.1330 5.4006
Indoor_2
(~100m)
0.2267 0.2043 0.1636
Indoor_3
(~100m)
1.5959 3.0011 6.8438
Outdoor_1
(~300m)
0.1378 1.9007 0.7280
Outdoor_2
(~400m)
0.9840 2.1944 1.2607
Outdoor_3
(~500m)
2.6659 4.7023 4.6066
@@
  • WHU-Helmet [30]: Collected on compact helmet system, the dataset contains complex large-scale scenes including wild, urban and tunnel, with challenges of drastic motion from wearer and repeated symmetrical geometric structures. The datasets used include Residence_5.2, Park_3.2, Subway_3.3, Mall_6.3.
@@
  • BotanicGarden [31]: Mounted on all-terrain wheeled robot, this dataset consists of diverse natural scenes including riversides, narrow trails and bridges. The datasets used include the Garden series.
@@
  • Handheld datasets: We collected multiple datasets of different scenarios and motion conditions, covering indoor and outdoor scenes like library, campus road and UAV airfield. The handheld data collection device is shown in Fig. 4. The datasets used include Indoor and Outdoor series.
In the evaluation of WHU-Helmet, Point-LIO diverged midway in the Residence_5.2 and Subway_3.3 due to severe motion combined with locally degraded scenarios, and diverged in Park_3.2 due to extensive occlusion by pedestrians. And in Mall_6.3, the system experienced significant drift during turns due to the reflection from glass walls of the shopping mall. In the first three datasets, our AC-LIO framework effectively resists the challenging scenarios mentioned above and outperformed the FAST-LIO in long-term, large-scale localization accuracy. Only in the last dataset of WHU-Helmet slightly lower than FAST-LIO.
For the BotanicGarden datasets, since the sensors are mounted on a wheeled robotic platform, the overall motion is smooth and suitable for evaluating the long-duration drift under regular motion conditions. Across the three LIO-rosbags provided by BotanicGarden, the AC-LIO demonstrates higher accuracy, exhibiting better convergence in long-term localization.
Next is the evaluation of handheld datasets. In Indoor_1 UAV airfield and Indoor_3 library scenarios, we conduct intense turning in narrow environment. And in the Indoor_2 garage scene, with smooth motion and highly structured surroundings, Point-LIO achieves higher closed-loop accuracy. Then in the Outdoor sequences, we handhold the device and attempt different motion conditions including running and sharp turns in campus road scenes.
Concrete experiment results are demonstrated in Table I, Table II and Fig. 5 . The results indicate the AC-LIO framework generally outperforms the FAST-LIO2 and Point-LIO in terms of accuracy during long-term, large-scale scenes, and exhibits high robustness under challenging scenarios.
@@
Figure 5: Accuracy evaluation in campus road scene. The figure shows the result of Outdoor_1. We keep running after initialization to enhance the motion intensity and complete a closed loop about 300m. The closure trajectory and the map consistency suggest that our AC-LIO achieves better performance.
IV-B Convergence Criteria Validity Evaluation
To evaluate the validity of the proposed convergence criteria, we examine the odometry accuracy at different $\eta$ threshold levels. While keeping other parameters the same, we only adjust the threshold of convergence criteria. Due to the factors such as dynamic objects, cumulative pointcloud map errors, etc., we fine-tune the threshold $\eta$ within the interval of $[\sigma_{t},2\sigma_{t}]$. The specific experimental results are shown in Table III.
  • $\eta=0.5\sigma_{t}$: Lower than the expected mean of pointcloud residuals, rarely perform backpropagation. (Only considering LiDAR ranging errors, $\mathbb{E}(\mathbf{z})=\sigma_{t}$)
TABLE III: Convergence Criteria Validity Evaluation - RMSE of ATE ($m$)
@@ IV-B Convergence Criteria Validity Evaluation
V CONCLUSIONS
In this paper, we propose AC-LIO, a novel asymptotically and consistently converging LIO framework. We reconsider the factors that may introduce errors in long-duration, large-scale scenarios, aiming to further enhance the convergence of the odometry. First, during iterative state estimation, we backpropagate the update term along the prior state chain and asymptotically compensate for residual distortion before the next iteration. Second, considering the weak correlation between the initial error and motion distortion in the current frame, we propose a convergence criteria based on pointcloud constraints to control backpropagation. In summary, the approach of guiding asymptotic compensation based on convergence criteria is designed to promote the consistent convergence of pointcloud registration, ensuring robustness while further improving the odometry accuracy.
We conducted a series of experiments on public and self-collected datasets, covering various scenarios and motion conditions across multiple scales. The experiments show that, compared to other state-of-the-art frameworks, our AC-LIO framework effectively promotes consistent convergence in state estimation and further enhances the accuracy of long-term, large-scale localization and mapping.

### v2

@@ III-D Convergence Criteria
Figure 3: The relationship between original ranging variance of LiDAR point and point-to-plane residual variance, where the residual $z$ is a scalar.
@@
To exploit the statistical properties of pointcloud, during the iterative state estimation, we select the mean value of the point-to-plane residual (abbreviated as APR above), as the criteria for evaluating the convergence of pointcloud registration. For simplification, assuming the pointcloud registration fully converges to the global optimum and only considering the ranging accuracy of LiDAR, we can roughly model the expectation of the APR and provide a threshold reference for the convergence criteria.
@@
Denote the incident angle of point $\mathbf{p}$ to the fitted plane $S$ as $\phi$, it can be expressed as $\cos^{-1}\tfrac{|\mathbf{u}\cdot\hat{\mathbf{p}}^{G}|}{|\mathbf{u}|\cdot|\hat{\mathbf{p}}^{G}|}$ according to Eq. 3. Note that $\mathbf{z}$ is the point-to-plane residual vector and $z$ is its scalar value. Assuming the ranging error of LiDAR point follows Gaussian distribution with zero mean and $\sigma$ variance. When the laser beam is perpendicular to the fitted plane, the residual $z_{\perp}$ follows chi-squared distribution and can be expressed as follows:
@@
$\begin{gathered}{z}_{\perp}=|\mathbf{p}^{G}-\hat{\mathbf{p}}^{G}|\sim\mathcal{N}(0,\sigma)\\ \frac{{z}_{\perp}^{2}}{\sigma^{2}}\sim\chi^{2}(1),\;\mathbb{E}({z}_{\perp})=\sigma\end{gathered}$ (10)
@@
As the surroundings vary during system’s ego-motion, the distribution of normal vector $\mathbf{u}$ to plane $S$ also changes. The variation of the geometrical scene would be reflected in the distribution of the incident angle $\phi$. From Fig. 3, the residual variance $\sigma_{z}$ should be $\sigma\cos\phi$. Assuming the incident angle is uniformly distributed over the domain $[0,\pi/2)$, then the expected mean of point-to-plane residual can be expressed as:
$\begin{gathered}\mathbb{E}({z})=\frac{2}{\pi}\int_{0}^{\frac{\pi}{2}}\mathbb{E}({z}_{\perp})\cos{\phi}\;d\phi=\frac{2}{\pi}\sigma\end{gathered}$ (11)
@@
Therefore, assuming the pointcloud registration converges to the global optimum, and only considering the LiDAR ranging error, the expectation of APR should be ${2\sigma}/{\pi}$. If the mean value of current point-to-plane residual is lower than this criteria, the last registration can be regarded as sufficiently converged.
@@
Based on this threshold of convergence criteria, we can easily determine the convergence degree of pointcloud registration, and control the backpropagation and asymptotic distortion compensation. Noting that $\sigma_{t}=\mathbb{E}({z})$ and the selected threshold of convergence criteria as $\eta$, due to the factors such as dynamic objects and cumulative errors in pointcloud map, the threshold $\eta$ can be appropriately extended to the interval of $[\sigma_{t},2\sigma_{t}]$ in practice.
@@
In detail, we perform two conditional judgments in iterative state estimation: First, evaluate the previous LiDAR frame, if its APR is less than the threshold $\eta$, it can be considered to be sufficiently converged and we will continue. Then, check the convergence of current registration, if the APR is still less than $\eta$, there is no need to backpropagate for small residual distortion; otherwise, we will perform the backpropagation and compensation during the iteration.
@@
On one hand, this approach can ensure the backpropagation and asymptotic compensation only performed as the previous LiDAR frame converged well, thus enhancing the odometry robustness in long-term and large-scale scenario. On the other hand, it can avoid lots of additional operations in unreliable or unnecessary conditions, to improve the system computational efficiency. The experiment results of our framework are shown in the next section.
@@ IV EXPERIMENT RESULTS
TABLE II: Odometry Accuracy Evaluation - RMSE of ATE & End to End Errors($m$)
Dataset LiDAR Type Sequence Ours FAST-LIO2 DLIO SLICT Point-LIO
WHU1 Helmet Livox AVIA Residence_5.2 (452.9m) 0.415 0.505 4.660 1.550 $\times$
Park_3.2 (628.9m) 0.582 0.618 7.725 0.602 $\times$
Subway_3.3 (827.0m) 1.670 3.752 3.688 1.967 $\times$
Mall_6.3 (1006.5m) 0.216 0.208 3.238 0.274 15.409
Botanic1 Garden Livox AVIA 1005_05 (566.8m) 0.773 1.053 0.640 7.931 0.917
1006_03 (686.2m) 1.796 2.075 3.017 6.527 2.471
1008_01 (855.9m) 1.152 1.577 14.829 4.445 1.590
VLP16 1005_05 (566.8m) 0.175 0.244 0.565 37.425 0.750
1006_03 (686.2m) 0.377 0.388 0.393 $\times$ 0.445
1008_01 (855.9m) 0.339 0.363 0.694 $\times$ 0.457
Handheld2 Livox AVIA Indoor_1 (~100m) 0.893 1.133 $\times$ 1.960 5.401
Indoor_2 (~100m) 0.227 0.204 3.354 0.249 0.164
Indoor_3 (~100m) 1.596 3.001 $\times$ 2.476 6.844
Outdoor_1 (~300m) 0.138 1.901 0.486 1.835 0.728
Outdoor_2 (~400m) 0.984 2.194 7.879 1.803 1.261
Outdoor_3 (~500m) 2.666 4.702 1.445 4.127 4.607
Average(RMSE) 0.750 1.078 3.945 7.590 3.148
  • 1  means the dataset is evaluated by RMSE of ATE, and  2 means it is evaluated by End to End errors.
  • $\times$  denotes the system severely diverged midway, and $\underline{\hskip 7.11317pt}$  denotes the second best result compared to bolded result.
In this section, we qualitatively and quantitatively evaluate AC-LIO framework through a series of experiments with various public and self-collected datasets. To examine the performance of our approach in long-term and large-scale scenarios, we conduct extensive experiments on the odometry accuracy, the time efficiency of system, and the validity of convergence criteria.
@@
IV-A Experiment Dataset
The datasets cover a wide range of structured environments (e.g., campus road and garage) and unstructured settings (e.g., botanical gardens), as well as different carriage modes (helmet, vehicle, handheld) to boost the challenge and completeness of assessment.
@@ IV-A Experiment Dataset
  • WHU-Helmet [40]: Collected on compact helmet system with Livox LiDAR, the dataset contains complex large-scale scenes including wild, urban and tunnel, with challenges of drastic motion from wearer and repeated symmetrical geometric structures.
@@
  • BotanicGarden [41]: Mounted on all-terrain wheeled robot, this dataset consists of diverse natural scenes including riversides, narrow trails and bridges with both Livox Avia and VLP16 in sequence.
@@
  • Handheld datasets: We collected multiple datasets of different scenarios and motion conditions, covering indoor and outdoor scenes like library, campus road and UAV airfield with Livox Avia. The handheld data collection device is shown in Fig. 4.
IV-B Odometry Accuracy Evaluation
@@
Figure 5: Accuracy evaluation in campus road scene. The figure shows the result of sequence Outdoor_1. Within the dashed box are, in order, the car, the streetlight, and the pattern on the ground. We keep running after initialization to enhance the motion intensity and complete a closed loop about 300m. The closure trajectory and the map consistency suggest that our AC-LIO achieves better performance.
In odometry accuracy experiments, we compare AC-LIO framework with current state-of-the-art LIO systems including FAST-LIO2, DLIO, SLICT and Point-LIO. Some other frameworks are designed for spinning LiDAR hence they’re not included in comparison. All algorithms are implemented in C++ and employed on an i9 CPU computer with the ROS in Ubuntu. These methods only utilize CPU for computation without the assistance of GPU parallel computing. For comparison, we try to fine-tune other methods’ parameters to achieve their best performance. Besides, we maintain the threshold of convergence criteria in the extended interval, i.e., $\eta\in[\sigma_{t},2\sigma_{t}]$.
In the evaluation of WHU-Helmet, the sequences Residence_5.2 and Subway_3.3 introduce aggressive motion combined with locally degraded scenarios, while Park_3.2 and Mall_6.3 suffer from extensive occlusion by pedestrians. These factors impose higher demands on the robustness of the LIO system. For the BotanicGarden datasets, we test both the solid-state (Livox AVIA) and the spinning (VLP16) LiDAR. Since the sensors are mounted on a wheeled robotic platform, the overall motion is smooth and suitable for evaluating long-duration drift under regular motion conditions. However, the abundance of unstructured scenes in the garden (e.g., thick woods, bushes) poses challenges for constructing residual constraints. In our self-collected datasets, we handheld the sensor suite and completed multiple closed-loop paths in different scenarios. In Indoor_1 UAV airfield and Indoor_3 library scenarios, we conduct intense turning in narrow environment. For the Outdoor sequences, we handhold the device and attempt different motion conditions including running and sharp turns in campus road scenes.
Concrete experiment results are demonstrated in Table II and Fig. 5. Our approach further optimizes the residual motion distortion via feedback mechanism based on the convergence criteria, thus improving the accuracy while ensuring the robustness of the system under challenging scenarios. The results indicate that our AC-LIO method generally outperforms other benchmark frameworks in terms of accuracy during long-term, large-scale scenes, with the average RMSE among sequences decreased by 30.4% compared to the second best, and exhibits high robustness under challenging scenarios.
TABLE III: Time Consumption Evaluation ($ms$)
Ours FAST-LIO2 DLIO SLICT Point-LIO
Min 4.65 3.15 3.43 21.33 11.95
Max 74.46 28.48 27.01 345.97 62.36
Mean 14.62 16.09 9.12 61.97 34.31
Median 11.71 16.68 8.50 51.49 34.98
Figure 6: The result of time efficiency evaluation. The cross line inside box denotes the median process time. Our AC-LIO achieves the second highest efficiency, only behind the lightweight DLIO framework. And the maximum iteration time also meets the real-time requirements.
IV-C Time Efficiency Evaluation
We calculate the runtime of each algorithm during iteration, with the results shown in Table III and Fig. 6. In terms of average and median processing time, the time efficiency of our AC-LIO framework ranks second place among the tested frameworks, just following the lightweight DLIO framework, and the maximum time for iteration also satisfies realtime requirement (below 100ms for 10Hz LiDAR input). Due to the convergence criteria strategy, our method can avoid incurring extra optimization costs in unreliable or unnecessary conditions, thus achieving the balance between system convergence and high time efficiency.
IV-D Convergence Criteria Validity Evaluation
TABLE IV: Convergence Criteria Validity Evaluation - RMSE of ATE ($m$)
@@ IV-D Convergence Criteria Validity Evaluation
To evaluate the validity of the proposed convergence criteria, we examine the odometry accuracy at different $\eta$ threshold levels. While keeping other parameters the same, we only adjust the threshold of convergence criteria. Due to the factors such as dynamic objects, cumulative pointcloud map errors, etc., we fine-tune the threshold $\eta$ within the interval of $[\sigma_{t},2\sigma_{t}]$. The specific experimental results are shown in Table IV.
  • $\eta=0.5\sigma_{t}$: Lower than the expected mean of pointcloud residuals, rarely perform backpropagation. (Only considering LiDAR ranging errors, $\mathbb{E}({z})=\sigma_{t}$)
V Conclusion and Future Work
In this paper, we proposed AC-LIO, a novel asymptotically and consistently converging LIO framework. We reconsider the factors that may introduce errors in long-duration, large-scale scenarios, aiming to further enhance the convergence of LIO system. The approach of guiding asymptotic compensation based on convergence criteria is designed to promote consistent convergence of pointcloud registration, further improving the odometry accuracy while ensuring system robustness. Numerous experiment results indicate that our framework’s average RMSE is reduced by about 30.4% over the second best result, achieving higher accuracy in long-term, large-scale localization and mapping.
Future Work: We notice the performance of back propagation depends on the guidance of convergence criteria to some extend. The introduction of convergence criteria is essentially an attempt to overcome the disruption of the initial error on the back propagation during the iteration update. In future work, we will continue to study the convergence assessment of odometry, the combination of asymptotic backpropagation with other strategies, and the perception and adaptation of degenerate scenarios, to further explore the error-analyzable and scenario-interpretable SLAM framework.